### v1

# Linearized Stability of Harada Thin-Shell Wormholes

Abstract
ABSTRACT
Using Darmois-Israel-Sen junction conditions, and with help of Visser’s cut-and-paste method, we study the dynamics of thin-shell wormholes that are made of two conformally Killing gravity (a.k.a Harada gravity) black holes. We check the energy conditions for different values of the new parameter that Harada introduced, as alternative for dark energy. We examine the radial acceleration to reveal the attractive and repulsive characteristics of the thin-shell wormhole throat. We consider the dynamics and stability of the wormhole around the static solutions of the linearized radial perturbations at the wormhole throat. Finally, we determine the regions of stability by applying the concavity test on the “speed of sound” as a function in the throat radius and other spacetime parameters, particularly the new Harada parameter.
pacs:
04.90.+e, 04.20.-q, 04.20.Gz
I Introduction
Traversable Lorentzian wormholes [1, 2] are known as solutions for Einstein’s field equations. The Conventional way to find a wormhole solution is to select some equations of state of matter such as Chaplygin gas [3], phantom energy [4, 5], and/or quintessence [6]. Then one can apply geometrical constructions such as thin-shell spacetimes [7], evolving wormholes [8], rotating spacetimes [9], dust shell wormholes [10], and/or Casimir wormholes [11] in order to alleviate the violation of energy conditions associated with the flaring-out condition that distinguishes wormhole solutions from other solutions. A thin-shell is a singular boundary hypersurface with energy-momentum conditions as stabilizers for such surface. If a thin shell is used to connect two separate spacetime with flaring-out condition, this thin shell forms a thin-shell wormhole. There are numerous studies that consider different black holes creating thin-shell wormholes in de-Sitter (dS) and anti-de-Sitter (AdS) spacetimes [12, 13, 14, 15, 16, 17]. Stability of these thin-shell wormholes are examined too [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37].
Recently, Harada targeted three main obstacles in Einstein’s general theory of relativity [38]. These obstacles are: i. the nature of the cosmological constant as an integration constant, ii. the derivation of conservation of energy-momentum tensor rather than being as conjectured an ad hoc, and iii. the usage of conformal flat metric as a vacuum solution and the unphysical solutions associated with it. Harada new gravity theory shows that accelerating expansion of the universe naturally appears as a consequence of the beyond general relativity gravitational field equations even in the absence of the conventional cosmological constant and/or dark energy [39]. Harada endeavors are considered a conformal Killing gravity theory in which the energy-momentum tensor is corrected by a divergenceless conformal Killing tensor, and thus the theory is compared to Cotton gravity [40]. Harada theory comes with many physical consequences such as generalized solutions of Schwarschild-Reissner-Nordström-AdS and regular black holes [41, 42], non-asymptotically flat traversable wormholes [43], black bounces [44], and pp-waves [45].
In this letter we study thin-shell wormholes in Harada gravity theory. In section ($\mathbf{II}$), we use Visser’s cut-and-paste technique [47, 46], together with Darmois-Israel-(Sen) junction conditions [48, 49, 50, 51], to connect two conformal Killing gravity regions of spacetime through a thin shell. The cut-and-paste method with the junction conditions are comprehensively reviewed in Ref. [52]. The methodology of cut-and-paste provides the advantages of the employment of lesser amount of exotic matter, and hence traveling objects through this wormhole avoid the regions occupied by exotic matter [47]. This results in confining the exotic matter at the thin-shell regions similar to the matter studied in other examples of thin-shell wormholes [20, 53, 36, 54, 55, 56, 57, 58, 59, 60, 61, 62]. Also, we study the components of the hypersurface energy-momentum tensor using the extrinsic curvature components. We utilize these components to obtain the stress and pressure then comment on the violation of energy conditions because of the exotic matter at the wormhole throat. From the discontinuity in the extrinsic curvature, we compare the effect of extra fourth order term in Harada metric on the spacetime to the effect of the usual second order cosmological constant term on spacetimes studied in Ref. [63]. Additionally we comment on the physics of attraction and repulsion on the wormhole throat in terms of the acceleration.
In section ($\mathbf{III}$), we analyze the linearized stability of Harada thin-shell wormhole by implementing the concavity test on the “speed of sound” as a function of spacetime parameters: the mass, the cosmological constant, and the new Harada parameter that corresponds to the fourth order term in the metric. Then, we visualize the change in stability regions upon varying the value of Harada parameter while we fix the values both mass and cosmological constant.
In section ($\mathbf{IV}$) we summarize the results of the previous two sections and comment on them.
II Dynamics of Thin-Shell Wormhole
The Harada black hole can be constructed [38] starting with a static spherically symmetric metric
$ds^{2}_{\text{Schw.}}=$ $-e^{\nu(r)}dt^{2}+e^{-\nu(r)}dr^{2}$ (1)
$+r^{2}d\theta^{2}+r^{2}\sin^{2}(\theta)d\phi^{2}\leavevmode\nobreak\ .$
The field equations are derived from the Harada totally symmetric tensor
$H_{\mu\nu\rho}=$ $\nabla_{\mu}R_{\nu\rho}+\nabla_{\nu}R_{\rho\mu}+\nabla_{\rho}R_{\mu\nu}$ (2)
$-\frac{1}{3}\left(g_{\nu\rho}\partial_{\mu}+g_{\rho\mu}\partial_{\nu}+g_{\mu\nu}\partial_{\rho}\right)R\leavevmode\nobreak\ ,$
where $R_{\mu\nu}$ is the Ricci tensor, and $R$ is the Ricci scalar. Harada tensor is subjected to the condition
$H_{\mu\nu\rho}=8\pi T_{\mu\nu\rho}\leavevmode\nobreak\ ,$ (3)
where the tensor condition is defined from the energy momentum tensor $T_{\mu\nu}$ and its trace $T$ as
$T_{\mu\nu\rho}=$ $\nabla_{\mu}T_{\nu\rho}+\nabla_{\nu}T_{\rho\mu}+\nabla_{\rho}T_{\mu\nu}$ (4)
$-\frac{1}{6}\left(g_{\nu\rho}\partial_{\mu}+g_{\rho\mu}\partial_{\nu}+g_{\mu\nu}\partial_{\rho}\right)T\leavevmode\nobreak\ .$
It is well established that Schwarzschild metric is obtained from defining $R_{11}$ and $R_{22}$ then plugging them in Einstein’s equations. Similarly, one can study $H_{111}$ together with other Ricci tensor components such that eq.(1) becomes:
$ds^{2}_{\text{H}}=-f(r)dt^{2}+f(r)^{-1}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}(\theta)d\phi^{2},$ (5)
$f(r)=1-\frac{2M}{r}-\frac{\Lambda}{3}r^{2}-\frac{N}{5}r^{4}\leavevmode\nobreak\ .$ (6)
Here $\Lambda$ is the cosmological constant, $M$ is the usual mass, and $N$ is the new Harada parameter that corresponds to implementing the conformal Killing conditions on the energy-momentum tensor. By finding the roots of the quintic polynomial  [^1] It is a textbook fact that Galois theory says there is no formula for the roots of quintic equation., i.e., $f(r)=0$ or
$3Nr^{5}+5\Lambda r^{3}-15r+30M=0\leavevmode\nobreak\ ,$ (7)
one can define the spatial regions where the inner, event ($r_{h}$) and cosmological ($r_{c}$) horizons are located. However, we need to avoid the combinations of $\Lambda$, $M$, and $N$ that lead to the formation of extreme black holes [15], where in such system the event and cosmological horizons coincide. This is crucial for preserving the throat radius $R$ of the wormhole as $r_{h}<R<r_{c}$.
Figure 1: The behavior of $g^{rr}$ metric component for (Harada: $M=1,\Lambda=10^{-2},N=10^{-4}$),
Reissner-Nordstrom de-Sitter
(RNdS: $M=1,\Lambda=10^{-2},Q=0.5$),
and Schwarzschild de-Sitter (SdS: $M=1,\Lambda=10^{-2}$).
Now we follow the cut-and-paste technique [46, 63] to construct a geodesically complete manifold ${\Gamma=\Gamma_{+}\bigcup\leavevmode\nobreak\ \Gamma_{-}}$. We start by cutting spacetime regions ${\Gamma_{\pm}:=\left\{r_{\pm}\leq R\leavevmode\nobreak\ |\leavevmode\nobreak\ R>r_{h}\right\}}$ inside the throat radius $R$. This is followed by pasting the timelike hypersurface regions, named a thin shell ${\partial\Gamma=\partial\Gamma_{+}\bigcup\leavevmode\nobreak\ \partial\Gamma_{-}}$ , where ${\partial\Gamma_{\pm}:=\left\{r_{\pm}=R\leavevmode\nobreak\ |\leavevmode\nobreak\ R>r_{h}\right\}}$, that bounds the bulk of two Harada black holes.
Next, we follow Darmois-Israel-(Sen) junction conditions [48, 50, 49] by defining the $\Gamma$ manifold coordinates as $x^{\mu}:=(t,r,\theta,\phi)$ and $\partial\Gamma$ shell coordinates as $\zeta^{i}:=(\tau,\theta,\phi)$, where $\tau$ is the proper time measured by a comoving observer when around the throat of the wormhole. The induced metric of the shell is given by
$ds^{2}_{\partial\Gamma}=-d\tau^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}(\theta)d\phi^{2},$ (8)
where the parametric equation $r=R(\tau)$ relates $\Gamma$ to $\partial\Gamma$. So in vacuum spacetime [64], the interior solution $r_{0}$ is matched to the exterior one $R$ at the junction surface $\partial\Gamma$ when the surface stress/pressure coefficients are present. We obtain surface stresses using the discontinuity in the extrinsic curvature $\mathcal{K}_{ij}$. Therefore, the thin-shell surface confines the exotic matter to a finite region [^2] See figure 1 of Ref. [64].. To minimize the violation of the average null energy condition (ANEC), we design the wormhole such that the exotic matter is impounded to the junction region $r_{0}<r<R$ with the limit $r_{0}\to R$ that causes the junction to evolve into a thin-shell.
After that, We decompose Harada spacetime using the Gauss-Codazzi approach, and consequently we obtain Israel’s junction condition on $\Gamma$ [52]. This condition is described by the $3D$ energy-momentum tensor on the junction $\mathcal{S}^{i}_{\leavevmode\nobreak\ j}=\text{diag}\left(-\sigma,p_{\theta},p_{\phi}\right)$ as
$\mathcal{S}_{ij}=-\frac{1}{8\pi}\left(\left[\mathcal{K}_{ij}\right]-\delta_{ij}\mathcal{K}\right),$ (9)
where $\left[\mathcal{K}_{ij}\right]=\mathcal{K}^{+}_{ij}-\mathcal{K}^{-}_{ij}$ is the discontinuity in the extrinsic curvature, and $\mathcal{K}=\left[\mathcal{K}^{i}_{\leavevmode\nobreak\ i}\right]$ is its trace.
We now define the unit vectors $n^{\pm}_{\mu}$ normal to $\partial\Gamma$ as
$n^{\pm}_{\mu}=\pm\left(\leavevmode\nobreak\ \biggl{\lvert}g^{\alpha\beta}\frac{\partial f}{\partial x^{\alpha}}\frac{\partial f}{\partial x^{\beta}}\biggr{\rvert}^{-1/2}\frac{\partial f}{\partial x^{\mu}}\right)\leavevmode\nobreak\ .$ (10)
Then, the components of the extrinsic curvature, or the second fundamental form, are defined as
$\mathcal{K}^{\pm}_{ij}=-n_{\mu}\left(\frac{\partial^{2}x^{\mu}}{\partial\zeta^{i}\zeta^{j}}+\Gamma^{\mu\pm}_{\nu\rho}\leavevmode\nobreak\ \frac{\partial x^{\nu}}{\partial\zeta^{i}}\frac{\partial x^{\rho}}{\partial\zeta^{j}}\right)$ (11)
We substitute eq.(6) in eq.(10) to get
$n^{\pm}_{\mu}=\biggl{(}\mp\dot{R},\pm\frac{\sqrt{\dot{R}^{2}+f(R)}}{f(R)},0,0\biggr{)}\leavevmode\nobreak\ .$ (12)
After that, we substitute eq.(12) in eq.(11) to get the extrinsic curvature components in Harada gravity as
$\mathcal{K}^{\pm}_{\theta\theta}=\mathcal{K}^{\pm}_{\phi\phi}=\pm\frac{1}{R}\sqrt{1-\frac{2M}{R}-\frac{\Lambda}{3}R^{2}-\frac{N}{5}R^{4}+\dot{R}^{2}}\leavevmode\nobreak\ ,$ (13)
$\mathcal{K}^{\pm}_{\tau\tau}=\mp\frac{1}{2}\frac{\frac{2M}{R^{2}}-\frac{2\Lambda}{3}R-\frac{4N}{5}R^{3}+2\ddot{R}}{\sqrt{1-\frac{2M}{R}-\frac{\Lambda}{3}R^{2}-\frac{N}{5}R^{4}+\dot{R}^{2}}}\leavevmode\nobreak\ .$
Thus we use these components to define the surface stress and pressure as
$\sigma$ $=-\frac{1}{2\pi}\mathcal{K}_{\theta\theta}$ (14)
$=-\frac{1}{2\pi R}\sqrt{1-\frac{2M}{R}-\frac{\Lambda}{3}R^{2}-\frac{N}{5}R^{4}+\dot{R}^{2}}\leavevmode\nobreak\ ,$
$p$ $=p_{\theta}=p_{\phi}=\frac{1}{4\pi}\left(\mathcal{K}_{\tau\tau}+\mathcal{K}_{\theta\theta}\right)$ (15)
$=\frac{1}{8\pi R}\bigg{(}\frac{\frac{2M}{R^{2}}-\frac{2\Lambda}{3}R-\frac{4N}{5}R^{3}+2\ddot{R}}{\sqrt{1-\frac{2M}{R}-\frac{\Lambda}{3}R^{2}-\frac{N}{5}R^{4}+\dot{R}^{2}}}R$
$\qquad\qquad+2\frac{1-\frac{2M}{R}-\frac{\Lambda}{3}R^{2}-\frac{N}{5}R^{4}+\dot{R}^{2}}{\sqrt{1-\frac{2M}{R}-\frac{\Lambda}{3}R^{2}-\frac{N}{5}R^{4}+\dot{R}^{2}}}\bigg{)}.$
And for the static configuration, i.e., $\dot{R}=\ddot{R}=0$, the surface stresses become
$\sigma_{0}=-\frac{1}{2\pi R_{0}}\sqrt{1-\frac{2M}{R_{0}}-\frac{\Lambda}{3}R_{0}^{2}-\frac{N}{5}R_{0}^{4}}\leavevmode\nobreak\ ,$ (16)
$p_{0}$ $=\frac{1}{8\pi R_{0}}\bigg{(}\frac{\frac{2M}{R_{0}^{2}}-\frac{2\Lambda}{3}R_{0}-\frac{4N}{5}R_{0}^{3}}{\sqrt{1-\frac{2M}{R_{0}}-\frac{\Lambda}{3}R_{0}^{2}-\frac{N}{5}R_{0}^{4}}}R_{0}$ (17)
$\qquad\qquad+2\frac{1-\frac{2M}{R_{0}}-\frac{\Lambda}{3}R_{0}^{2}-\frac{N}{5}R_{0}^{4}}{\sqrt{1-\frac{2M}{R_{0}}-\frac{\Lambda}{3}R_{0}^{2}-\frac{N}{5}R_{0}^{4}}}\bigg{)}\leavevmode\nobreak\ .$
The last two equations are used to study the violation of different energy conditions; the surface density $\sigma_{0}<0$ imposes the violation of the weak energy condition (WEC), meanwhile the null energy condition (NEC), $\sigma_{0}+p_{0}>0$, is maintained as long as ${f(R_{0})<\frac{M}{R_{0}}-\frac{\Lambda}{3}R_{0}^{2}-\frac{2N}{5}R_{0}^{4}}$ with no additional exotic effects. And for the strong energy condition (SEC), $\sigma_{0}+3p_{0}>0$, it is also maintained as long as ${f(R_{0})>\Lambda R_{0}^{2}+\frac{6N}{5}R_{0}^{4}-\frac{3M}{R_{0}}}$.
(a) $N=10^{-4}$.
(b) $N=10^{-6}$.
Figure 2: The energy conditions expressed in terms of $\sigma$ and $p$ vs. the throat radius $R_{0}$ with fixed $M=1$ and $\Lambda=10^{-2}$, and different values of Harada coefficient: $N=10^{-4}$ for figure.2.(a) and $N=10^{-6}$ for figure.2.(b).
For Harada black hole with no radial pressure, $p_{r}=0$, and a mass density localized at the throat $\rho=\sigma_{0}\leavevmode\nobreak\ \delta(r-R_{0})$, the total amount of exotic matter necessary to keep the wormhole open is
[…]
$=-2R_{0}\sqrt{1-\frac{2M}{R_{0}}-\frac{\Lambda}{3}R_{0}^{2}-\frac{N}{5}R_{0}^{4}}\leavevmode\nobreak\ .$
Thus we can study the attraction and repulsion characteristics of the corresponding thin-shell wormhole by examining the four-acceleration $a^{\mu}=u^{\nu}\nabla_{\nu}u^{\mu}$, where the four-velocity $u^{\mu}=(1/\sqrt{f(r)},0,0,0)$ has only time component ${u^{\mu}\equiv dt/d\tau}$. The geodesic equation of a test particle defines the acceleration as
$\frac{d^{2}r}{d\tau^{2}}=-a^{r}\leavevmode\nobreak\ ,$ (19)
where in Harada gravity it is given by
$a^{r}=\Gamma^{r}_{tt}\left(\frac{dt}{d\tau}\right)^{2}=\frac{6N}{15}r^{3}+\frac{\Lambda}{3}r-\frac{M}{r^{2}}\leavevmode\nobreak\ .$ (20)
We notice that the wormhole has attractive or repulsive nature if $a^{r}>0$ or $a^{r}<0$ respectively.
(a) Different attractive and repulsive behaviors of $a^{r}$ at small $R$ values for different $N$ values.
(b) Convergent behavior of $a^{r}$ at small $R$ values for different $N$ values. $a^{r}$ becomes repulsive at almost the same value of $R$ regardless the value of $N$.
Figure 3: Attraction and repulsion in terms of acceleration $a^{r}$ vs. the throat radius $R$ with fixed $M=1$ and $\Lambda=10^{-2}$, and different values of Harada coefficient $N$.
III Linearized Stability Analysis
We check the stability of the wormhole by performing linear perturbation for eq.(16) and eq.(17) around the static configuration [63] ,i.e., when $(R=R_{0})$. Then, we differentiate eq.(14) with respect to $\tau$ to obtain the continuity equation
$\frac{d(\sigma A)}{d\tau}+p\frac{dA}{d\tau}=0\leavevmode\nobreak\ ,$ (21)
which directly yields
$\sigma^{\prime}=-\frac{2}{R}(\sigma+p)\leavevmode\nobreak\ ,$ (22)
where $A=4\pi R^{2}$ is the area of the wormhole throat, and $\sigma^{\prime}=\dot{\sigma}/\dot{R}$; the dot is for $d/d\tau$, and the prime is for $d/dR$.
Next, we rearrange eq.(14) and define a potential function
$V(R)=f(R)-4\pi^{2}R^{2}\sigma^{2}=-\dot{R}\leavevmode\nobreak\ .$ (23)
Then, we substitute eq.(22) in the first derivative of eq.(23) to find that
$V^{\prime}(R)=\frac{2M}{R^{2}}-\frac{2\Lambda}{3}R-\frac{4N}{5}R^{3}+8\pi^{2}R\sigma(\sigma+2p)\leavevmode\nobreak\ .$ (24)
And for the second derivative of (23), pressure is parameterized to be a function in the density $p:=p(\sigma)$ [7]. Thus we introduce another parameter $\vartheta(\sigma)=dp/d\sigma$, which can be seen as the “speed of sound”. Therefore, the second derivative of (23) becomes
$V^{\prime\prime}(R)$ $=f^{\prime\prime}(R)-8\pi^{2}\left[2\sigma(\sigma+p)(1+2\vartheta)+(\sigma+2p)^{2}\right]$ (25)
$=f^{\prime\prime}(R)+\Bigg{[}\leavevmode\nobreak\ \leavevmode\nobreak\ \frac{1}{R^{2}}\bigg{(}Rf^{\prime}(R)-2f(R)\bigg{)}\bigg{(}1+2\vartheta\bigg{)}$
$\qquad\qquad\quad\leavevmode\nobreak\ \leavevmode\nobreak\ -\frac{1}{2}\frac{\left(f^{\prime}(R)\right)^{2}}{f(R)}\Bigg{]}\leavevmode\nobreak\ .$
In order to linearize the time rate change of the radius $\dot{R}$, we utilize Taylor expansion with the potential function around the static point $R=R_{0}$
$V(R)=$ $V(R_{0})+(R-R_{0})V^{\prime}(R_{0})$ (26)
$+\frac{1}{2}(R-R_{0})^{2}V^{\prime\prime}(R_{0})+\mathcal{O}\left[(R-R_{0})^{3}\right]\leavevmode\nobreak\ .$
We use eq.(16) and eq.(17) to evaluate eq.(23) and eq.(24) at $R=R_{0}$. So we notice that the first two terms in the expansion vanish $V(R_{0})=V^{\prime}(R_{0})=0$. Meanwhile the second derivative term eq.(25) becomes
$V^{\prime\prime}(R_{0})=$ $-\frac{4M}{R_{0}^{3}}-\frac{2\Lambda}{3}-\frac{12N}{5}R_{0}^{2}$ (27)
$+\bigg{[}\frac{2M}{R_{0}^{3}}-\frac{2\Lambda}{3}-\frac{4N}{5}R_{0}^{2}$
$-\frac{2}{R_{0}}+\frac{4M}{R_{0}^{2}}+\frac{2\Lambda}{3}R_{0}+\frac{2N}{5}R_{0}^{3}\bigg{]}\bigg{[}1+2\vartheta_{0}\bigg{]}$
$-\frac{1}{2}\frac{\left(\frac{2M}{R_{0}^{2}}-\frac{2\Lambda}{3}R_{0}-\frac{4N}{5}R_{0}^{3}\right)^{2}}{1-\frac{2M}{R_{0}}-\frac{\Lambda}{3}R^{2}_{0}-\frac{N}{5}R_{0}^{4}}\leavevmode\nobreak\ .$
On can use $(1+2\vartheta)=(\sigma^{\prime}+2p^{\prime})/\sigma^{\prime}$ to express $\vartheta$ in terms of the metric parameters $M,\leavevmode\nobreak\ N$ and $\Lambda$ such that $V^{\prime\prime}$ has no implicit dependency of the metric parameters $\vartheta(\Lambda,M,N)$. But our current focus is to study the behavior of $\vartheta$ when the throat is stable.
The concave down test $V^{\prime\prime}(R_{0})<0$ results in causing either expansion or contraction of the throat when small perturbation takes place. While the convex, or the concave up, condition $V^{\prime\prime}(R_{0})>0$ stabilizes the throat with a local minimum of $V(R_{0})$. Therefore, we solve eq.(27) for $\vartheta_{0}$ at that local minimum to get
$\vartheta_{0}<\frac{1}{2}\bigg{[}\frac{\frac{\left(\frac{2M}{R_{0}^{2}}-\frac{2\Lambda}{3}R_{0}-\frac{4N}{5}R_{0}^{3}\right)^{2}}{2(1-\frac{2M}{R_{0}}-\frac{\Lambda}{3}R^{2}_{0}-\frac{N}{5}R_{0}^{4})}+\frac{4M}{R_{0}^{3}}+\frac{2\Lambda}{3}+\frac{12N}{5}R_{0}^{2}}{\frac{2M}{R_{0}^{3}}-\frac{2\Lambda}{3}-\frac{4N}{5}R_{0}^{2}-\frac{2}{R_{0}}+\frac{4M}{R_{0}^{2}}+\frac{2\Lambda}{3}R_{0}+\frac{2N}{5}R_{0}^{3}}-1\bigg{]}\leavevmode\nobreak\ .$ (28)
(a) $N=0.001$.
(b) $N=0.01$.
(c) $N=1$.
(d) $N=10$.
Figure 4: Regions of stability of the thin-shell wormhole for the Bardeen de-Sitter solution for fixed values of $M=1$ and $\Lambda=10^{-2}$, and different values of $N$. Stable regions are the blue shaded domains.
IV Discussion
In this letter we design Harada thin-shell wormholes. We utilize Visser’s technique of cut-and-paste with Darmois-Israel-(Sen) junction condition to connect two Harada regions of spacetime through a thin shell. We compare the asymptotic behavior of Harada metric with that of SdS and RNdS as in fig.(1). Also, We examine the components of the energy-momentum hypersurface tensor using the second fundamental form. Next, we find that WEC is always violated. However, both NEC and SEC can be maintained upon imposing the inequalities that relate $f(r)$ to $f^{\prime}(r)$. The energy conditions are shown in fig.(2). Then, we study the radial acceleration to examine the attractive and repulsive nature of the wormhole throat. The results are plotted in fig.(3).
After that, we analyze the linearized stability of Harada thin-shell wormhole by examining the concavity behavior on the “speed of sound” as a function in Harada parameter, the mass, and the cosmological constant. And we notice the change in stability regions upon varying Harada parameter while both mass and cosmological constant are constant. The analysis is demonstrated in fig.(4). We conclude that for a lesser value of cosmological constant and lesser/larger value of magnetic charge, relative to the amount of mass, regions of stability vary. Once the mass is equal to the Harada parameter, we no longer have stability regions. Therefore in the Harada spacetime, and for a small value of the cosmological constant, thin-shell wormholes are kept open as long as the value of Harada parameter is away from the value of mass.
V Acknowledgement
The authors would like to thank the Department of Physics, Santa Clara University, especially the Chair Prof. Dr. G. Ramon, for supporting the senior undergraduate physics seminars where the authors met and worked to produce this letter.

### v2

@@
 Stability of these thin-shell wormholes are examined too [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37].
+Also, one should consider several other examples of wormhole solutions that do not require exotic matter such as those supported by the Dirac-Maxwell fields [38, 39] and those supported by applying different geometrical modifications to the usual spacetime [40, 41, 42, 43, 44, 45].
+It is worth noting that the stability of wormholes is an old issue in theories of gravity, and until this moment there is no wormhole solution for which stability has been realized under general perturbations, see the examples in Ref.
+[46, 47, 48] and the cited references there.
 Recently, Harada targeted three main obstacles in Einstein’s general theory of relativity [49].
@@
 Harada endeavors are considered a conformal Killing gravity theory in which the energy-momentum tensor is corrected by a divergenceless conformal Killing tensor, and thus the theory is compared to Cotton gravity [51].
-Harada theory comes with many physical consequences such as generalized solutions of Schwarschild-Reissner-Nordström-AdS and regular black holes [41, 42], non-asymptotically flat traversable wormholes [43], black bounces [44], and pp-waves [45].
-In this letter we study thin-shell wormholes in Harada gravity theory.
+Harada theory comes with many physical consequences [52, 53] such as generalized solutions of Schwarschild-Reissner-Nordström-AdS and regular black holes [54, 55], non-asymptotically flat traversable wormholes [56], black bounces [57], and pp-waves [58].
+We are interested in studying the stability of thin-shell wormholes in Harada gravity theory for the following two main reasons.
+The first one is that finding a stable soltuion to traversable wormholes in conformal Killing gravity is a prominent result because such solution is not allowed by the null energy condition in the context of the conventional general theory of relativity [59].
+The second reason is that there is a class of wormhole solutions of conformal Killing gravity found in Ref.
+[56] the does not demand the support of exotic matter or dark energy to stabilize the wormholes; and this stimulates considering the thin-shell wormholes within the same theory.
 In section ($\mathbf{II}$), we use Visser’s cut-and-paste technique [61, 60], together with Darmois-Israel-(Sen) junction conditions [62, 63, 64, 65], to connect two conformal Killing gravity regions of spacetime through a thin shell.
@@
 the throat radius $R_{0}$ with fixed $M=1$ and $\Lambda=10^{-2}$, and different values of Harada coefficient: $N=10^{-4}$ for figure.2.(a) and $N=10^{-6}$ for figure.2.(b).
-For Harada black hole with no radial pressure, $p_{r}=0$, and a mass density localized at the throat $\rho=\sigma_{0}\leavevmode\nobreak\ \delta(r-R_{0})$, the total amount of exotic matter necessary to keep the wormhole open is
+For Harada black hole with no radial pressure and a mass density localized at the throat $\rho=\sigma_{0}\leavevmode\nobreak\ \delta(r-R_{0})$, the total amount of exotic matter necessary to keep the wormhole open is
 […]
@@
 where $A=4\pi R^{2}$ is the area of the wormhole throat, and $\sigma^{\prime}=\dot{\sigma}/\dot{R}$; the dot is for $d/d\tau$, and the prime is for $d/dR$.
-Next, we rearrange eq.(14) and define a potential function
+Next, rearrange eq.(14) and define a potential function
 $V(R)=f(R)-4\pi^{2}R^{2}\sigma^{2}=-\dot{R}\leavevmode\nobreak\ .$ (23)
@@
 (d) $N=10$.
-Figure 4: Regions of stability of the thin-shell wormhole for the Bardeen de-Sitter solution for fixed values of $M=1$ and $\Lambda=10^{-2}$, and different values of $N$.
+Figure 4: Regions of stability of the thin-shell wormhole for the Harada solution with fixed values of $M=1$ and $\Lambda=10^{-2}$, and different values of $N$.
 Stable regions are the blue shaded domains.
@@
 The results are plotted in fig.(3).
-After that, we analyze the linearized stability of Harada thin-shell wormhole by examining the concavity behavior on the “speed of sound” as a function in Harada parameter, the mass, and the cosmological constant.
-And we notice the change in stability regions upon varying Harada parameter while both mass and cosmological constant are constant.
-The analysis is demonstrated in fig.(4).
-We conclude that for a lesser value of cosmological constant and lesser/larger value of magnetic charge, relative to the amount of mass, regions of stability vary.
+After that, we analyze the linearized stability of Harada thin-shell wormhole by examining the concavity behavior of eq.(28), that of the “speed of sound”, as a function in Harada parameter, the mass, and the cosmological constant.
+And we notice changes in stability regions upon varying Harada parameter while keeping the mass and cosmological constant fixed.
+Besides being considered as a stability measure for the wormhole solutions, the speed of sound function in principle can help detecting such systems by revealing the thermodynamics of the equations of states of the ordinary and/or exotic matter supporting the wormhole system.
+The analysis of stability regions of speed of sound is demonstrated in fig.(4), and we conclude that for a lesser value of cosmological constant and lesser/larger value of magnetic charge, relative to the amount of mass, regions of stability vary.
 Once the mass is equal to the Harada parameter, we no longer have stability regions.